a Mer sehne Mer Senne pr me otler then3
Answer #1

3 is not a quadratic residue modulo 3 as there exists no solution for the equation

r 3 mod 3 as

12 13 mod 3 22-43 mod 3

So we consider the case when

p2

Now \frac{q-1}{2}=\frac{2^p-2}{2}=2^{p-1}-1

Also by Law of Quadratic Reciporocity

3 (q-1)3-1

Hence

(p-1)(3-1 3

Since p is odd so

2p = 22H 142-1

Now

4\equiv 1\mod 3\\ \implies 4^r.2\equiv 2 \mod 3

Thus

q= 4^r.2-1\equiv 1 \mod 3

Thus q is a quadratic residue modulo 3.

Hence

(\frac{q}{3})=1\\

Using (*) we have

(\frac{q}{3})=-1

Thus q is a quadratic non-residue modulo 3